# Interpretable global minima of deep ReLU neural networks on sequentially separable data

### v2

@@ 5. Hyperplanes and sequentially separable data
$H(p,\nu):=\left\{p+x\in\mathbb{R}^{n}:\langle x,\nu\rangle=0\right\}.$ (5.1)
For data that is sequentially linearly separable, we have the corresponding hyperplanes $H_{j}=H(p_{j},h_{j})$ for some unit vectors $h_{j}$, $j=1,\cdots,Q$. We will always consider data sets to be finite, and so the distance from any such hyperplane to the each of the two sets it is separating is always positive. Moreover, a pair $(p,\nu)\subset\mathbb{R}^{n}\times\mathbb{R}^{n}$, along with an angle $\theta<\pi$, define a pair of cones $p+\mathfrak{C}_{\theta}[\nu]$ and $p+\mathfrak{C}_{\theta}[-\nu]$ as in (3.18), and so to each hyperplane $H(p,\nu)$ is associated a family of cones.
Theorem 5.2.
@@
$\min_{(W_{i},b_{i})_{i=1}^{L}}\mathcal{C}[(W_{i},b_{i})_{i=1}^{L}]=0,$ (5.2)
with weights and biases given recursively by (3.4), (3.5) for the following cumulative parameters: for $\ell=1,\cdots,Q$,
@@
$p_{j}+\mathfrak{C}_{\theta_{j}}[h_{j}]$ $\supset\bigcup_{i<j}\{p_{i}\}\cup\bigcup_{i>j}\mathcal{X}_{0,i}$
and (5.6)
@@
$\bigcup_{j>1}\mathcal{X}_{0,j}$ $\subset p_{1}+\mathfrak{C}_{\theta_{1}}[h_{1}].$ (5.7)
@@
$W_{1}=W^{(1)}:=W_{\theta_{1}}R_{1},$ (5.8)
for $W_{\theta_{1}}$ defined as in (3.23), and $R_{1}\in O(d_{0},\mathbb{R})$ such that
$R_{1}h_{1}=\frac{u_{d_{0}}}{\left\lvert{u_{d_{0}}}\right\rvert},$ (5.9)
@@
$b_{1}=b^{(1)}:=-W_{1}p_{1}.$ (5.10)
Then Lemma 3.4 implies we have the following truncation map
$\tau_{W^{(1)},b^{(1)}}(x)=\begin{cases}x,\quad&x\in p_{1}+\mathfrak{C}_{\theta_{1}}[h_{1}],\\ p_{1},\quad&x\in p_{1}+\mathfrak{C}_{\theta_{1}}[-h_{1}],\end{cases}$ (5.11)
@@
$X_{0,j}^{(\tau,1)}=\tau_{W^{(1)},b^{(1)}}(X_{0,j})=\begin{cases}p_{1}u_{N_{1}}^{T},\quad&j=1,\\ X_{0,j},\quad&j>1,\end{cases}$ (5.12)
@@
$X_{0,j}^{(\tau,\,\ell-1)}=\begin{cases}p_{j}u_{N_{j}}^{T},\quad&j\leq\ell-1,\\ X_{0,j},\quad&j>\ell-1,\end{cases}$ (5.13)
where $p_{j}$ are the points given by (5.5). We will construct cumulative parameters $W^{(\ell)},b^{(\ell)}$ such that
$X_{0,j}^{(\tau,\ell)}=\tau_{W^{(\ell)},b^{(\ell)}}(X_{0,j}^{(\tau,\ell-1)})=\begin{cases}p_{j}u_{N_{j}}^{T}=X_{0,j}^{(\tau,\ell-1)},\quad&j<\ell,\\ p_{\ell}u_{N_{\ell}}^{T},\quad&j=\ell,\\ X_{0,j}=X_{0,j}^{(\tau,\ell-1)},\quad&j>\ell.\end{cases}$ (5.14)
By sequential linear separability of the data, there exists a hyperplane $H_{\ell}=H(p_{\ell},h_{\ell})$ in the collection given by (5.5) separating $\mathcal{X}_{0,\ell}$ and $\bigcup_{j<\ell}\{p_{j}\}\cup\bigcup_{j>\ell}\mathcal{X}_{0,j}$. We can use the induction hypothesis (5.13) to rewrite
$\mathcal{X}_{0,\ell}=\mathcal{X}_{0,\ell}^{(\tau,\ell-1)}\,\text{and }\,\bigcup_{j<\ell}\{p_{j}\}\cup\bigcup_{j>\ell}\mathcal{X}_{0,j}=\bigcup_{j\neq\ell}\mathcal{X}_{0,j}^{(\tau,\ell-1)},$ (5.15)
@@
$\bigcup_{j\neq\ell}\mathcal{X}_{0,j}^{(\tau,\ell-1)}\subset p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[h_{\ell}]\quad\text{ and }\quad\mathcal{X}_{0,\ell}\subset p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[-h_{\ell}].$ (5.16)
@@
$W^{(\ell)}:=W_{\theta_{\ell}}R_{\ell},$
$b^{(\ell)}:=-W^{(\ell)}\,p_{\ell},$ (5.17)
for $W_{\theta_{\ell}}$ defined as in (3.23) and $R_{\ell}\in O(n,\mathbb{R})$ such that
$R_{\ell}h_{\ell}=\frac{u_{d_{0}}}{\left\lvert{u_{d_{0}}}\right\rvert},$ (5.18)
so that Lemma 3.4 gives
$\tau_{W^{(\ell)},b^{(\ell)}}(x)=\begin{cases}x,\quad&x\in p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[h_{\ell}],\\ p_{\ell},\quad&x\in p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[-h_{\ell}],\end{cases}$ (5.19)
which, along with (5.16), implies (5.14).
@@
$X^{(Q+1)}$ $=W^{(Q+1)}(X_{0})^{(\tau,Q)}+B^{(Q+1)},$ (5.20)
and we wish to minimize
$\min_{W^{(Q+1)},b^{(Q+1)}}\left\lVert{W^{(Q+1)}(X_{0})^{(\tau,Q)}+B^{(Q+1)}-Y^{ext}}\right\rVert^{2}_{\mathcal{L}^{2}}.$ (5.21)
@@
$W^{(Q+1)}(\overline{X_{0}^{red}})^{(\tau,Q)}=Y.$ (5.25)
As the collection $\{p_{j}\}_{j=1}^{Q}$ is linearly independent,
$(\overline{X_{0}^{red}})^{(\tau,Q)}=\left[p_{1}\cdots p_{Q}\right]$ (5.26)
@@
$((\overline{X_{0}^{red}})^{(\tau,Q)}){}^{+}(\overline{X_{0}^{red}})^{(\tau,Q)}={\textbf{1}}_{Q\times Q},$ (5.27)
@@
is a solution to (5.25), and we have in fact produced a collection of weights and biases $(W_{i}^{**},b_{i}^{**})_{i=1}^{Q+1}$ given by the cumulative parameters defined in (5) and (5.28) that achieves zero cost,
$\mathcal{C}[(W_{i}^{**},b_{i}^{**})_{i=1}^{Q+1}]=0.$ (5.29)
@@
Finally, the minimum we have constructed is degenerate: for a fixed collection of hyperplanes, the angles $(\theta_{\ell})_{\ell=1}^{Q}$ of the associated cones can vary in $(\theta_{1,min},\pi)\times\cdots\times(\theta_{Q,min},\pi)$ (recall (5)), and so parametrize a family of cumulative weights $(W^{(\ell)}[\theta_{\ell}])_{\ell=1}^{Q}$. ∎
6. Proof of Proposition 4.1

### v3

@@ 5 Hyperplanes and Sequentially Separable Data
$H(p,\nu):=\left\{p+x\in\mathbb{R}^{n}:\langle x,\nu\rangle=0\right\}.$
For data that is sequentially linearly separable, we have the corresponding hyperplanes $H_{j}=H(p_{j},h_{j})$ for some unit vectors $h_{j}$, $j=1,\cdots,Q$. We will always consider data sets to be finite, and so the distance from any such hyperplane to the each of the two sets it is separating is always positive. Moreover, a pair $(p,\nu)\subset\mathbb{R}^{n}\times\mathbb{R}^{n}$, along with an angle $\theta<\pi$, define a pair of cones $p+\mathfrak{C}_{\theta}[\nu]$ and $p+\mathfrak{C}_{\theta}[-\nu]$ as in (12), and so to each hyperplane $H(p,\nu)$ is associated a family of cones.
Remark 4.
Similarly, we note that there exists $\epsilon>0$ such that each point $p_{j}$ can be freely perturbed on an $\epsilon$-ball around it (possibly moving the corresponding hyperplane $H_{j}$ as well) so that $(\tilde{p}_{j},\tilde{H}_{j})_{j=1}^{Q}$ still satisfies Definition 5, for $\tilde{p}_{j}\in(p_{j}+B_{\epsilon}(p_{j}))\cap\tilde{H}_{j}$, $j=1,\cdots,Q$. Therefore, we may assume that $\{p_{j}\}_{j=1}^{Q}$ form a linearly independent set.
Theorem 6.
@@
$\min_{(W_{i},b_{i})_{i=1}^{L}}\mathcal{C}[(W_{i},b_{i})_{i=1}^{L}]=0,$
with weights and biases given recursively by (5), (6) for the following cumulative parameters: for $\ell=1,\cdots,Q$,
@@
$p_{j}+\mathfrak{C}_{\theta_{j}}[h_{j}]$ $\supset\bigcup_{k<j}\{p_{k}\}\cup\bigcup_{k>j}\mathcal{X}_{0,k}$
and (22)
@@
$\bigcup_{j>1}\mathcal{X}_{0,j}$ $\subset p_{1}+\mathfrak{C}_{\theta_{1}}[h_{1}].$ (23)
@@
$W_{1}=W^{(1)}:=W_{\theta_{1}}R_{1}\in\mathbb{R}^{d_{0}\times d_{0}},$
for $W_{\theta_{1}}$ defined as in (16), and $R_{1}\in O(d_{0},\mathbb{R})$ such that
$R_{1}h_{1}=\frac{u_{d_{0}}}{\left\lvert{u_{d_{0}}}\right\rvert},$
@@
$b_{1}=b^{(1)}:=-W_{1}p_{1}.$
Then Lemma 2 implies we have the following truncation map
$\tau_{W^{(1)},b^{(1)}}(x)=\begin{cases}x,\quad&x\in p_{1}+\mathfrak{C}_{\theta_{1}}[h_{1}],\\ p_{1},\quad&x\in p_{1}+\mathfrak{C}_{\theta_{1}}[-h_{1}],\end{cases}$
@@
$X_{0,j}^{(\tau,1)}=\tau_{W^{(1)},b^{(1)}}(X_{0,j})=\begin{cases}p_{1}u_{N_{1}}^{T},\quad&j=1,\\ X_{0,j},\quad&j>1,\end{cases}$
@@
$X_{0,j}^{(\tau,\,\ell-1)}=\begin{cases}p_{j}u_{N_{j}}^{T},\quad&j\leq\ell-1,\\ X_{0,j},\quad&j>\ell-1,\end{cases}$ (24)
where $p_{j}$ are the points given by (21). We will construct cumulative parameters $W^{(\ell)},b^{(\ell)}$ such that
$X_{0,j}^{(\tau,\ell)}=\tau_{W^{(\ell)},b^{(\ell)}}(X_{0,j}^{(\tau,\ell-1)})=\begin{cases}p_{j}u_{N_{j}}^{T}=X_{0,j}^{(\tau,\ell-1)},\quad&j<\ell,\\ p_{\ell}u_{N_{\ell}}^{T},\quad&j=\ell,\\ X_{0,j}=X_{0,j}^{(\tau,\ell-1)},\quad&j>\ell.\end{cases}$ (25)
By sequential linear separability of the data, there exists a hyperplane $H_{\ell}=H(p_{\ell},h_{\ell})$ in the collection given by (21) separating $\mathcal{X}_{0,\ell}$ and $\bigcup_{j<\ell}\{p_{j}\}\cup\bigcup_{j>\ell}\mathcal{X}_{0,j}$. We can use the induction hypothesis (24) to rewrite
$\mathcal{X}_{0,\ell}=\mathcal{X}_{0,\ell}^{(\tau,\ell-1)}\,\text{and }\,\bigcup_{j<\ell}\{p_{j}\}\cup\bigcup_{j>\ell}\mathcal{X}_{0,j}=\bigcup_{j\neq\ell}\mathcal{X}_{0,j}^{(\tau,\ell-1)},$
@@
$\bigcup_{j\neq\ell}\mathcal{X}_{0,j}^{(\tau,\ell-1)}\subset p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[h_{\ell}]\quad\text{ and }\quad\mathcal{X}_{0,\ell}\subset p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[-h_{\ell}].$ (26)
@@
$W^{(\ell)}:=W_{\theta_{\ell}}R_{\ell}\in\mathbb{R}^{d_{0}\times d_{0}},$
$b^{(\ell)}:=-W^{(\ell)}\,p_{\ell}\in\mathbb{R}^{d_{0}},$ (27)
for $W_{\theta_{\ell}}$ defined as in (16) and $R_{\ell}\in O(d_{0},\mathbb{R})$ such that
$R_{\ell}h_{\ell}=\frac{u_{d_{0}}}{\left\lvert{u_{d_{0}}}\right\rvert},$
so that Lemma 2 gives
$\tau_{W^{(\ell)},b^{(\ell)}}(x)=\begin{cases}x,\quad&x\in p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[h_{\ell}],\\ p_{\ell},\quad&x\in p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[-h_{\ell}],\end{cases}$
which, along with (26), implies (25).
@@
$X^{(Q+1)}$ $=W^{(Q+1)}(X_{0})^{(\tau,Q)}+B^{(Q+1)},$
and we wish to minimize  [^6] Recall the matrices defined in (2).
$\min_{W^{(Q+1)},b^{(Q+1)}}\left\lVert{W^{(Q+1)}(X_{0})^{(\tau,Q)}+B^{(Q+1)}-Y^{ext}}\right\rVert^{2}_{\mathcal{L}^{2}}.$
@@
$W^{(Q+1)}(\overline{X_{0}^{red}})^{(\tau,Q)}=Y.$ (28)
As the collection $\{p_{j}\}_{j=1}^{Q}$ is assumed linearly independent (see Remark 4),
$(\overline{X_{0}^{red}})^{(\tau,Q)}=\left[p_{1}\cdots p_{Q}\right]$
@@
$((\overline{X_{0}^{red}})^{(\tau,Q)}){}^{+}(\overline{X_{0}^{red}})^{(\tau,Q)}={\textbf{1}}_{Q\times Q},$
@@
is a solution to (28), and we have in fact produced a collection of weights and biases $(W_{i}^{**},b_{i}^{**})_{i=1}^{Q+1}$ given by the cumulative parameters defined in (5) and (29) that achieves zero cost,
$\mathcal{C}[(W_{i}^{**},b_{i}^{**})_{i=1}^{Q+1}]=0.$
@@
Finally, the minimum we have constructed is degenerate: for a fixed collection of hyperplanes, the angles $(\theta_{\ell})_{\ell=1}^{Q}$ of the associated cones can vary in $(\theta_{1,min},\pi)\times\cdots\times(\theta_{Q,min},\pi)$, and so parametrize a family of cumulative weights $(W^{(\ell)}[\theta_{\ell}])_{\ell=1}^{Q}$.
6 Additional Remarks
We make here a few remarks concerning generalization and neural collapse.
6.1 Generalization
Let $\rho\in\mathbb{P}(Z)$ be a probability distribution on $Z=X\times Y$, the set of data $x\in X$ and labels $y\in Y\subset\mathbb{R}$. Following the most general setup in the book by Vapnik (2013), we consider
$\left\{Q(z,\alpha),\alpha\in\Lambda\right\}$
to be a parametrized set of functions representing the composition of a loss function with a function $f_{\alpha}:X\to Y$. The real risk of a function $f_{\alpha}$ on data distributed according to $\rho$ is
$R(\alpha):=\int_{Z}Q(z,\alpha)\,\text{{d}}\rho(z).$
Given a training set $\left\{z_{1},\cdots,z_{n}\right\}$, the empirical risk is
$R_{emp}(\alpha):=\frac{1}{n}\sum_{i=1}^{n}Q(z_{i},\alpha).$
Then if $a\leq Q(z,\alpha)\leq b$, the following holds for any $\alpha\in\Lambda$ and with probability at least $1-\eta$:
$R(\alpha)\leq R_{emp}(\alpha)+(b-a)\sqrt{\frac{VC(\Lambda)\log(n+1)-\log(\frac{\eta}{4})}{n}},$
where $VC(\Lambda)$ is the Vapnik–Chervonenkis dimension of the function class parametrized by $\Lambda$. In particular, for an empirical risk minimizer which achieves zero loss,
$R(\alpha_{*})\leq(b-a)\sqrt{\frac{VC(\Lambda)\log(n+1)-\log(\frac{\eta}{4})}{n}}.$ (30)
In this work, $X=\mathbb{R}^{d_{0}}$, $Y=\mathbb{R}^{Q}$, the functions $f_{\alpha}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{Q}$ are ReLU neural networks parametrized by weights and biases, and we consider the squared loss, so
$Q(z=(x,y),\alpha)=\left\lvert{f_{\alpha}(x)-y}\right\rvert^{2}.$
The loss function can be bounded, for instance, by imposing bounds on the norms of the parameters and the input data. See work by Chen et al. (2025) for a more detailed discussion.
If the VC dimension of the function class is finite, [^7] It is known that the VC dimension of real valued ReLU neural networks is finite, and explicit bounds exist (see e.g., Bartlett et al., 2019). then (30) implies that the results obtained in the work at hand generalize well to test data which is identically distributed to the training data, as they do not depend on the size of the training data set.
6.2 Cost Decomposition and Neural Collapse
Note that since the constructions in this paper do not depend on the cost function chosen, we can change it to average over each class separately, [^8] Equivalently, we could consider balanced classes.
$=\left\lVert{X^{(L)}-Y^{ext}}\right\rVert^{2}_{\mathcal{L}_{\mathcal{N}}^{2}},$
where $\mathcal{N}:=\operatorname{diag}(N_{j}{\textbf{1}}_{N_{j}\times N_{j}}|j=1,\cdots,Q)$ and $\langle A,B\rangle_{\mathcal{L}_{\mathcal{N}}^{2}}:=\operatorname{Tr}(A\mathcal{N}^{-1}B^{T})$. For simplicity, throughout this section we will refer to $\mathcal{C}_{\mathcal{N}}$ as $\mathcal{C}$. This can be decomposed as follows
$=\sum_{j=1}^{Q}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left\lvert{x_{0,j,i}^{(L)}-\overline{x_{0,j}^{(L)}}}\right\rvert^{2}+\left\lvert{\overline{x_{0,j}^{(L)}}-y_{j}}\right\rvert^{2}\right)$ (31)
$=\sum_{j=1}^{Q}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left\lvert{W^{(L)}\Delta x_{0,j,i}^{(\tau,L-1)}}\right\rvert^{2}+\sum_{j=1}^{Q}\left\lvert{W^{(L)}\overline{x_{0,j}^{(\tau,L-1)}}+b^{(L)}-y_{j}}\right\rvert^{2}.$
Since both terms are non-negative, the cost will be zero if, and only if, each term is zero, and so we have the following interpretation for the results presented in this paper: the truncation maps implemented by the hidden layers have the role of minimizing the first term, by which all variances are reduced to zero, while the last (affine) layer minimizes the second term, whereby the class averages are matched with the reference outputs $y_{j}$.
Han, Papyan, and Donoho (2022) observed that when training a classifier to zero mean square loss, the following phenomena are observed on the penultimate layer features:
  • (NC1) Within-class variability collapse, which for instance occurs if $\Delta X_{0}^{(L-1)}\to 0$.
  • (NC2) Convergence to simplex ETF: the vectors $(\overline{x_{0,j}^{(L-1)}}-\overline{x^{(L-1)}})$ evolve to have maximal angle between them, and the same length, forming an equilateral simplex, or simplex equiangular tight frame (ETF).
  • (NC3) Convergence to self-duality: $f_{j}^{(L-1)}+w_{j}\to 0$, where $f_{j}^{(L-1)}$ are unit vectors pointing from $\overline{x_{0,j}^{(L-1)}}$ to $\overline{x^{(L-1)}}$, defined analogously to (19), and $w_{j}$ are the rows of $W_{L}$, for $j=1,\cdots,Q$.
They consider balanced classes and decompose the mean square loss into the sum of a least-squares term and the remainder, and then further decompose the least-squares term into terms associated to the properties above:
$\mathcal{C}$ $=\mathcal{C}_{LS}+\mathcal{C}_{LS}^{\perp}=(\mathcal{C}_{NC1}+\mathcal{C}_{NC2,3})+\mathcal{C}_{LS}^{\perp}.$ (32)
Further, they provide empirical evidence that $\mathcal{C}_{LS}^{\perp}$ becomes negligible fast during training.
We will make some remarks comparing (32) with the cost decomposition (6.2) and the results in this paper. The last line of (6.2) can be written in two different ways,
$\mathcal{C}$ $=\sum_{j=1}^{Q}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left\lvert{W^{(L)}\Delta x_{0,j,i}^{(\tau,L-1)}}\right\rvert^{2}+\sum_{j=1}^{Q}\left\lvert{W^{(L)}\overline{x_{0,j}^{(\tau,L-1)}}+b^{(L)}-y_{j}}\right\rvert^{2}$ (33)
$=\sum_{j=1}^{Q}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left\lvert{W_{L}\Delta x_{0,j,i}^{(L-1)}}\right\rvert^{2}\,\,+\,\,\,\sum_{j=1}^{Q}\left\lvert{W_{L}\overline{x_{0,j}^{(L-1)}}+b_{L}-y_{j}}\right\rvert^{2}.$ (34)
If we let $b_{L}=0$ and $W_{L}=W_{LS}$ be the least squares matrix that minimizes the second term in (34), then $\mathcal{C}=\mathcal{C}_{LS}$, the first term in (34) is $\mathcal{C}_{NC1}$, and the second term is $\mathcal{C}_{NC2,3}$ (up to constants). Comparing with the interpretation given after (6.2), this suggests that the truncation maps in the hidden layers drive NC1, while the last-layer affine map is responsible for NC2,3.
Let us compare the last-layer weights and biases obtained by minimizing according to (34) and (33). For the rest of this section we take $Y={\textbf{1}}_{Q\times Q}$, $\overline{X^{(\tau,\ell)}}:=\left[\overline{x_{0,1}^{(\tau,\ell)}}\cdots\overline{x_{0,Q}^{(\tau,\ell)}}\right]$ and $\overline{X^{(\ell)}}:=\left[\overline{x_{0,1}^{(\ell)}}\cdots\overline{x_{0,Q}^{(\ell)}}\right]$. In this work we have minimized with respect to (33),
$\tilde{W}^{(L)}=(\overline{X^{(\tau,L-1)}}){}^{+},\,b^{(L)}=0,$
which yields
$\tilde{W}_{L}$ $=\tilde{W}^{(L)}(W^{(L-1)}){}^{+}$
$=(\overline{X^{(\tau,L-1)}}){}^{+}(W^{(L-1)}){}^{+}.$
Minimizing with respect to (34), as Han et al. (2022), gives
$W_{L}$ $=(\overline{X^{(L-1)}}){}^{+},\,b_{L}=0.$
We rewrite $W_{L}$: from Proposition 2,
$\overline{X^{(L-1)}}$ $=W^{(L-1)}\overline{X^{(\tau,L-1)}}+B^{(L-1)}$
$=W^{(L-1)}\left(\overline{X^{(\tau,L-1)}}+\beta^{(L-1)}\right)$
for
$\beta^{(L-1)}:=(W^{(L-1)}){}^{+}B^{(L-1)},$
assuming $W^{(L-1)}$ is surjective. Define the orthogonal projectors
$\mathcal{P}_{W}$ $:=\hskip 18.00005pt\mathcal{P}_{(\ker W^{(L-1)})^{\perp}}$ $=(W^{(L-1)}){}^{+}W^{(L-1)},$ (35)
$\mathcal{P}_{X}$ $:=\mathcal{P}_{\operatorname{ran\,}(\overline{X^{(\tau,L-1)}}+\beta^{(L-1)})}$ $=\left(\overline{X^{(\tau,L-1)}}+\beta^{(L-1)}\right)\left(\overline{X^{(\tau,L-1)}}+\beta^{(L-1)}\right)^{+}.$ (36)
Then
$W_{L}$ $=\left(W^{(L-1)}\left(\overline{X^{(\tau,L-1)}}+\beta^{(L-1)}\right)\right)^{+}$
=$\stackrel{{\scriptstyle(1)}}{{=}}\left(\mathcal{P}_{W}\left[\overline{X^{(\tau,L-1)}}+\beta^{(L-1)}\right]\right)^{+}\left(W^{(L-1)}\mathcal{P}_{X}\right)^{+},$
where $(1)$ follows from a result that can be found in the book by Campbell and Meyer (2009, Theorem 1.4.1). Now, observe that
$\operatorname{ran\,}\left(\overline{X^{(\tau,L-1)}}+\beta^{(L-1)}\right)\subset\operatorname{ran\,}(W^{(L-1)}){}^{+}=\left(\ker W^{(L-1)}\right)^{\perp},$ (37)
so
$\mathcal{P}_{W}\mathcal{P}_{X}=\mathcal{P}_{X}\mathcal{P}_{W}=\mathcal{P}_{X},$
and we conclude
$W_{L}$ $=\left(\overline{X^{(\tau,L-1)}}+\beta^{(L-1)}\right)^{+}\left(W^{(L-1)}\mathcal{P}_{X}\right)^{+}.$
If the sets in (37) are equal, then $\mathcal{P}_{W}=\mathcal{P}_{X}$ and $W_{L}$ further simplifies to
$W_{L}$ $=\left(\overline{X^{(\tau,L-1)}}+\beta^{(L-1)}\right)^{+}\left(W^{(L-1)}\right)^{+},$
so $W_{L}$ and $\tilde{W}_{L}$ would differ only in the $\beta^{(L-1)}$ term. This could be achieved, for instance, by taking $d_{L-1}=Q$ and $\overline{X^{(\tau,L-1)}}\in\mathbb{R}^{M\times Q}$ injective (recall we are already assuming $W^{(L-1)}$ to be surjective). Suppose this is the case, for simplicity, and consider two subcases:
  • First, if $b^{(L-1)}=0$, then
    $W_{L}=\tilde{W}_{L}\text{ and }b^{(L)}=W_{L}b^{(L-1)}+b_{L}=b_{L}=0$
    and we are in the setting of Han et al. (2022). This constrains the cone at the last nonlinear layer to be based at the origin.
  • If we wish to consider strong neural collapse (SNC) as defined by Mixon et al. (2022), then  [^9] Note that SNC is not reasonable if the penultimate layer features come from a ReLU neural network, as $\overline{x^{(L-1)}}=0$ would imply that $x_{0,j,i}^{(L-1)}=0$ for all $j=1,\cdots,Q$, $i=1,\cdots,N_{j}$. Still, the study of the ranks of the matrices and comparison with the first subcase is instructive.
    $\overline{x^{(L-1)}}=Q^{-1}\sum_{j=1}^{Q}\overline{x_{0,j}^{(L-1)}}=0,$
    which implies that $\overline{X^{(L-1)}}\in\mathbb{R}^{Q\times Q}$ cannot be full rank. To conserve injectivity of $\overline{X^{(\tau,L-1)}}$, we must have $b^{(L-1)}\neq 0$, and therefore $\tilde{W}_{L}\neq W_{L}$. Indeed, SNC assumes
    $W_{L}W_{L}^{T}=\sqrt{Q}\left({\textbf{1}}_{Q\times Q}-\frac{1}{Q}u_{Q}u_{Q}^{T}\right),\quad b_{L}=\frac{1}{Q}u_{Q},$
    and in that case not only is the last-layer weight matrix not the least squares matrix (6.2), it is not surjective. To fit this, the constructions in Proposition 3 and Theorem 6 could be modified so that $W^{(L)}\in\mathbb{R}^{Q\times M}$ has rank $Q-1$ while still achieving zero loss:
    $W^{(L)}=\left({\textbf{1}}_{Q\times Q}-B^{(L)}\right)\overline{X^{(\tau,L-1)}}^{+},\quad b^{(L)}=\frac{1}{Q}u_{Q}.$
The remarks in this section indicate that the constructions presented in this paper align well with the framework of neural collapse, at least in certain cases or with minor adaptations. Therefore, employing the truncation map as a tool to investigate the analytic and geometric properties of deep neural networks could also be advantageous in this context.
7 Proof of Proposition 3
To prove the proposition, we use barycentric coordinates. We consider a point $x\in\mathbb{R}^{M}$ to be represented by
$x=\sum_{j=1}^{Q}\kappa_{j}\overline{x_{0,j}}+\tilde{x}$ (38)
where $\tilde{x}\in(\operatorname{span}\{\overline{x_{0,j}}\}_{j=1}^{Q})^{\perp}$, and the means $\overline{x_{0,j}}$ form the vertices of a $Q$-simplex for which $\overline{x}$ is the barycenter (see Figure 1). Switching to these coordinates can be done when preprocessing the data.